Pumped-storage hydroelectricity allows energy to be stored to prepare for periods of high demand. Let a high-level reservoir be 500 m above a low-level reservoir. What is the useful power output if the water from the high-level reservoir flows at a rate of 100 kg per minute and the turbine is 85% efficient?


1
Expert's answer
2022-01-31T11:25:00-0500

Explanations & Calculation


  • Energy received by the generator during a period of a minute is

"\qquad\qquad
\begin{aligned}
\small E&=\small mgh \\

\end{aligned}"

  • Then the received power, which is the energy per unit time is

"\qquad\qquad
\begin{aligned}
\small P_r&=\small \frac{mgh}{t}=\frac{m}{t}.gh\\
\end{aligned}"

  • Finally, the power output is

"\qquad\qquad
\begin{aligned}
\small P_{out}&=\small P_r\times\text{efficiency}\\
\small P_{out}&=\small \big(\frac{m}{t}\Big).gh\times\frac{85}{100}
\end{aligned}"


  • You can give it a try substituting the given values in the formula above. m/t is the rate of water flow: 100kg per minute.
  • Per-minute needs to be converted into per second.